30 percent of households say they would feel secure if they had​ $50,000 in savings. You randomly select 8 households and ask them if they would feel secure if they had​ $50,000 in savings. Find the probability that the number that say they would feel secure is​ (a) exactly​ five, (b) more than​ five, and​ (c) at most five.
Found 2 solutions by stanbon, Boreal:
Answer by stanbon(75887)

30 percent of households say they would feel secure if they had​ $50,000 in savings. You randomly select 8 households and ask them if they would feel secure if they had​ $50,000 in savings.
Find the probability that the number that say they would feel secure is​
(a) exactly​ five:: p(x = 5) = 8C5*0.3^5*0.7^3 = binompdf(8,0.3,5) = 0.0467
(b) more than​ five::P(6<=x<=8) = 1 - binomcdf(8,0.3,5) = 0.0113
(c) at most five::P(0<=x<=5) = binomcdf(8,0.3,5) = 0.9887
-------------
Cheers,
Stan H.
----------------

Answer by Boreal(15235)

exactly 5: 8C5(0.3)^5(0.7)^3=66*0.000833=0.055 (rounded at end)
6:8C6(0.3)^6(0.7)^2=0.010
7:=8(0.3)^7)(0.7)=0.001
8:0.3^8=0.00007
Greater than 5, the last three: 0.011
At most 5 is everything else, the complement of greater than 5, and that is 0.989.